Mathematics 241 – Multivariable Calculus Quiz 6 – October 29, 2010 Your Name: Let f(x, y) = cos(x + y) + sin(x − y). A) (10) Find the directional derivative D_{(\sqrt{3}/2,1/2)}f(π/2, π/2).

B) (10) Find the tangent plane to z = f(x, y) at (π/2, π/2, −1).